1. On the first day of the fiscal year, a company issues a $4,400,000, 8%, 7-year...

1.

On the first day of the fiscal year, a company issues a $4,400,000, 8%, 7-year bond that pays semiannual interest of $176,000 ($4,400,000 × 8% × ½), receiving cash of $3,767,084.

Journalize the first interest payment and the amortization of the related bond discount. Round to the nearest dollar. If an amount box does not require an entry, leave it blank.

2.

On the first day of the fiscal year, a company issues a $4,300,000, 10%, 9-year bond that pays semiannual interest of $215,000 ($4,300,000 × 10% × ½), receiving cash of $4,561,441.

Journalize the first interest payment and the amortization of the related bond premium. Round to the nearest dollar. If an amount box does not require an entry, leave it blank.

3.

A $910,000 bond issue on which there is an unamortized premium of $83,000 is redeemed for $812,000.

Journalize the redemption of the bonds. If an amount box does not require an entry, leave it blank.

4.

Effect of Financing on Earnings per Share

Domanico Co., which produces and sells biking equipment, is financed as follows:

Bonds payable, 10% (issued at face amount) $900,000
Preferred $2 stock, $20 par 900,000
Common stock, $25 par 900,000

Income tax is estimated at 40% of income.

Determine the earnings per share on common stock, assuming that the income before bond interest and income tax is (a) $297,000, (b) $387,000, and (c) $477,000.

Enter answers in dollars and cents, rounding to two decimal places.

a. Earnings per share on common stock $

b. Earnings per share on common stock $

c. Earnings per share on common stock $
Answer #1
1
Interest expense 221208
       Discount on Bonds payable 45208 =(4400000-3767084)/7*6/12
       Cash 176000
2
Interest expense 200475
Premium on Bonds payable 14525 =(4561441-4300000)/9*6/12
       Cash 215000
3
Bonds payable 910000
Premium on Bonds payable 83000
      Cash 812000
      Gain on bond redemption 181000

4

a
Income before bond interest and income tax 297000
Less: Bond interest expense 90000
Income before income tax 207000
Less: Income tax expense 40% 82800
Net income 124200
Less: Preferred Dividends 90000 =(900000/20)*2
Net income for Common Stockholders 34200
Divide by Common Shares outstanding 36000 =900000/25
Earnings per share on common stock 0.95
b
Income before bond interest and income tax 387000
Less: Bond interest expense 90000
Income before income tax 297000
Less: Income tax expense 40% 118800
Net income 178200
Less: Preferred Dividends 90000
Net income for Common Stockholders 88200
Divide by Common Shares outstanding 36000
Earnings per share on common stock 2.45
c
Income before bond interest and income tax 477000
Less: Bond interest expense 90000
Income before income tax 387000
Less: Income tax expense 40% 154800
Net income 232200
Less: Preferred Dividends 90000
Net income for Common Stockholders 142200
Divide by Common Shares outstanding 36000
Earnings per share on common stock 3.95